### v1

@@ 5.4 Integrals of motion dependent on $x_{3}$
Let us present the generic form of operators $\tilde{Q}_{2}$ and $\tilde{Q}_{3}$ whose commutator with $P_{3}$ is proportional to $Q_{2}$ and $Q_{3}$ (97) respectively:
$\begin{split}&\tilde{Q}_{2}=\frac{\omega}{2}(\{P_{1},L_{2}\}+\{P_{2},L_{1}\})+\kappa(P_{1}L_{1}-P_{2}L_{2})\\ &+\frac{1}{2}(\lambda^{32}\{P_{3},L_{1}\}-\lambda^{31}\{P_{3},L_{2}\})+\left(N_{3}\cdot H\right)+\eta_{3}\end{split},$ (130)
$\tilde{Q}_{3}=\tilde{Q}_{2}-\mu L_{3}D-\nu\{L_{2},L_{3}\}-\frac{\alpha}{2}\{P_{3},K_{3}\}$ (131)
In accordance with (129) operators $Q_{2},Q_{3}$ and $\tilde{Q}_{2},Q_{3}$ satisfy the following relation
$[P_{3},\tilde{Q}_{\alpha}]={\mathrm{i}}Q_{c},c=2,3$ (132)
where $Q_{2}$ is operator (86) and $Q_{3}$ is operator (97) with $\sigma=0$. Notice that equation (132) has no solutions for $c=3,\sigma\neq 0$.
Let operator $\tilde{Q}_{c}$ with some combinations of arbitrary parameters $\omega,\kappa,\lambda^{31},\lambda^{32},\mu,\nu,\alpha$ be an integral of motion of equation (1) with arbitrary elements (9). Then operator $Q_{c}$ with the same values of the mentioned parameters is the integrals of motion too. Thus the necessary condition of the commutativity $\tilde{Q}_{c}$ with the Hamiltonian $H$ is the commutativity of $Q_{c}$ with $H$. Since we already found such operators $Q_{c}$, we know all admissible combinations of arbitrary parameters in integrals of motion (130). However, the mentioned necessary condition is not sufficient one, and so we are supposed to verify wether operators $\tilde{Q}_{c}$ commute with the found Hamiltonians. Making this routine job for the admissible values of arbitrary parameters fixed in (99)-(101) we recover the $x_{3}$ dependent integrals of motion presented in Items 2, 5 of Table 3 and Items 1-4 of Table 4.
A bit more sophisticated speculations are requested to find the integral of motion $\hat{Q}=\{P_{2},L_{1}\}+2\left(\frac{c_{1}x_{3}}{x_{2}^{2}}\cdot H\right)-2\frac{c_{2}x_{3}}{x_{2}^{2}}$ presented in Item 6 of Table 2. It satisfied the following relation:
@@
$\begin{split}&\hat{Q}_{1}=L_{1}^{2}-\left(\frac{c_{1}x_{3}^{2}}{x_{2}^{2}}\cdot H\right)+\frac{c_{2}x_{3}^{2}}{x_{2}^{2}},\\ &\hat{Q}_{2}=\{L_{2},L_{3}\}-\frac{2c_{6}x_{3}\tilde{r}+c_{4}x_{3}(x_{1}^{2}+2x_{2}^{2})}{x_{1}^{2}\tilde{r}}\\ &+\left(\frac{2c_{3}x_{3}\tilde{r}+c_{1}x_{3}(x_{1}^{2}+2x_{2}^{2})}{x_{1}^{2}\tilde{r}}\cdot H\right)\end{split}$ (136)
@@
first of which is presented in Items 6 of Table 3 and Item 2 of Table 5 while the second one can be found in Item 3 of Table 5.
@@
Notice that integrals of motion whose differential part $Q_{(n)}$ (36) being third and second order polynomials in $x_{3}$ are not admitted by the considered shift invariant PDM systems.
@@ 6 Discussion
We classify inequivalent quantum mechanical systems with position dependent masses which admit second order integrals of motion and one out of two one parametric Lie groups, i.e., dilatation and shift ones. The total number of such systems is equal to twenty seven, including nine ones invariant w.r.t. the dilatation transformations and eighteen shift invariant systems. The classification results are presented in five tables and include eleven integrable, ten superintegrable and six maximally superintegrable systems. Let us note that this statement is a bit conventional since the obtained systems are defined up to arbitrary functions and (or) arbitrary parameters, and in fact we enumerate the subclasses of the systems under study including the infinite sets of them. One more note is that the systems represented in Item 5 of Table 1 and Item 8 of Table 3 form the subclasses of the systems represented in Item 4 of Table 1 and Item 7 of Table 3 correspondingly.
The presented results can be treated as the next step in our programm of the complete classification of the 3d PDM systems admitting second order integrals of motion. Such systems possessing the symmetry w.r.t. three and two parametric Lie groups were classified in paper [42] and [45] respectively. The next step presupposes the classification of the systems admitting at least one parametric Lie symmetry group. In accordance with [24], up to equivalence it is sufficient to restrict ourselves to the two subclassess of the groups. The first of them includes rotations around the fixed axis, shifts along this axis and dilatations, we call these symmetries natural. The second subclass includes the so called combined transformations and includes superpositions of shifts and rotations, dilatations and rotations, and shift, rotations and conformal transformations. The main result of the present paper consists in completing the classification of the systems admitting the natural symmetries started in paper [47].
Thus to complete the classification of the 3d PDM systems admitting second order integrals of motion and a one parametric Lie symmetry group it is sufficient to classify the the systems admitting the combined symmetries mentioned in the above. This work is in progress.

### v3

@@ 4.2.4 Integrals of motion dependent on $x_{3}$
Let us present the generic form of operators $\tilde{Q}^{(1)}$ and $\tilde{Q}^{(2)}$ whose commutator with $P_{3}$ is proportional to $Q^{(1)}$ (96) and $Q^{(2)}$ (97) respectively:
$\tilde{Q}^{(1)}=\frac{\omega}{2}(\{P_{1},L_{2}\}+\{P_{2},L_{1}\})+\kappa(P_{1}L_{1}-P_{2}L_{2})+\nu\{L_{2},L_{3}\}+\left(N^{(1)}\cdot H\right)+\tilde{N}^{(1)},$ (174)
$\tilde{Q}^{(2)}=\mu L_{3}D+\frac{\alpha}{2}\{P_{3},K_{3}\}+\frac{1}{2}(\lambda^{32}\{P_{3},L_{1}\}-\lambda^{31}\{P_{3},L_{2}\})+\left(N^{(2)}\cdot H\right)+\tilde{N}^{(2)}$ (175)
In accordance with (173) operators $Q^{(1)},Q^{(2)}$ and $\tilde{Q}^{(1)},\tilde{Q}^{(2)}$ satisfy the following relation
$[P_{3},\tilde{Q}^{(a)}]={\mathrm{i}}Q^{(a)},a=1,2$ (176)
where $Q^{(1)}$ is operator (96) and $Q^{(2)}$ is operator (97) with $\sigma=0$. Notice that equation (176) has no solutions for $a=2,\sigma\neq 0$.
Let operator $\tilde{Q}^{(a)}$ with some combinations of arbitrary parameters be an integral of motion of equation (1) with arbitrary elements (6). Then operator $Q^{(a)}$ with the same values of the mentioned parameters is the integral of motion too. Thus the necessary condition of the commutativity $\tilde{Q}^{(a)}$ with the Hamiltonian $H$ is the commutativity of $Q^{(a)}$ with $H$. Since we already found all inequivalent operators $Q^{(a)}$, we know all admissible combinations of arbitrary parameters in integrals of motion (174). However, the mentioned necessary condition is not sufficient one, and so we are supposed to verify whether operators $\tilde{Q}^{(a)}$ do commute with the found Hamiltonians. Making this routine job we recover the $x_{3}$ dependent integrals of motion presented in Items 2, 5 of Table 2 and Items 1-4 of Table 3.
A bit more sophisticated speculations are requested to find the integral of motion $\hat{Q}=\{P_{2},L_{1}\}+2\left(\frac{c_{1}x_{3}}{x_{2}^{2}}\cdot H\right)-2\frac{c_{2}x_{3}}{x_{2}^{2}}$ presented implicitly in Item 3 of Table 3 as a commutator of the first integral of motion with $P_{3}$. It satisfies the following relation:
@@
$\begin{split}&\hat{Q}_{1}=L_{1}^{2}-\left(\frac{c_{1}x_{3}^{2}}{x_{2}^{2}}\cdot H\right)+\frac{c_{2}x_{3}^{2}}{x_{2}^{2}},\\ &\hat{Q}_{2}=D^{2}-c_{1}(r^{2}\cdot H)+c_{2}r^{2},\\ &\hat{Q}_{3}=\{L_{2},L_{3}\}-\frac{2c_{6}x_{3}\tilde{r}+c_{4}x_{3}(x_{1}^{2}+2x_{2}^{2})}{x_{1}^{2}\tilde{r}}\\ &+\left(\frac{2c_{3}x_{3}\tilde{r}+c_{1}x_{3}(x_{1}^{2}+2x_{2}^{2})}{x_{1}^{2}\tilde{r}}\cdot H\right)\end{split}$ (180)
@@
first of which is presented in Item 3 of Table 3 and Item 2 of Table 4 while the second and third ones one can be found in Item 6 of Table 3 and Item 3 of Table 4.
@@
Notice that integrals of motion whose differential part $Q_{(n)}$ (22) includes third and second order polynomials in $x_{3}$ are not admitted by the considered shift invariant PDM systems.
4.3 Classification tables
We have completed the classification of the PDM systems which are invariant with respect to rotations around the third coordinate axis and admit second order integrals of motion. In this section we represent the obtained results in the three tables which include integrable, superintegrable and maximally superintegrable systems.
In the tables we give the inverse masses and potentials of such systems together with the functionally independent integrals of motion. Notice that the number of the the linearly independent integrals is more extended. To find them it is sufficient to calculate the commutators of the presented operators with $P_{3}$.
In Table 2 we use the following notations: $\Omega_{m}={(\tilde{\omega}+\tilde{r}^{2})^{m}+(-1)^{m+1}(\tilde{\omega}-\tilde{r}^{2})^{m}},$ $y=x_{1}^{2}-x_{2}^{2}-1,\ \tilde{\omega}=\sqrt{\tilde{r}^{4}-4y},c_{m}$ with $m=1,2,..$ are arbitrary parameters. In addition, in this and the following system notation (43) is used.
Table 2 represents all inequivalent PDM systems which admit two integrals of motion, one of which is the generator of shift transformations along the third coordinate axis and the other is a second order integral of motion. In other words the systems presented in Table 2 are integrable while superintegrable systems are represented in Tables 3 and 4.
The potentials and inverse masses presented in Table 2 include arbitrary functions of some special variables. In other words the class of such systems is rather extended.
The systems represented in Item 8 are particular cases of the generic system given in Item 7. The masses and potentials of these systems are polynomials in variables $\tilde{\omega}\pm\tilde{r}^{2}$ while in Item 7 there are arbitrary functions of these variables. The nice feature of the polynomial solutions is that the related integrals of motion can be presented explicitly while for the case of arbitrary functions we are able to represent the constructive elements $N$ and $\tilde{N}$ of these integrals of motion only as solutions of the related equations (124) which can be solved for any particular case of these arbitrary functions.
No $f$ $V$ Integrals of motion
2 $\frac{\tilde{r}^{2}}{F(\tilde{r})-G(\varphi)}$ $\frac{\tilde{F}(\varphi)+\tilde{G}(\tilde{r})}{G(\varphi)-F(\tilde{r})}$ $L_{3}^{2}+(G(\varphi)\cdot H)+\tilde{G}(\varphi)$
4 $\frac{\tilde{r}^{2}}{c_{1}\tilde{r}^{2}+F(\varphi)}$ $\frac{c_{2}\tilde{r}^{2}+G(\varphi)}{c_{1}\tilde{r}^{2}+F(\varphi)}$$\{P_{3},D\}-2c_{1}(x_{3}\cdot H)+2c_{2}x_{3}$
6 $\frac{\tilde{r}^{2}}{c_{1}\tilde{r}^{2}+F(\nu\varphi-\alpha\ln(\tilde{r}))}$ $\frac{c_{2}\tilde{r}^{2}+G(\nu\varphi-\alpha\ln(\tilde{r}))}{c_{1}\tilde{r}^{2}+F(\nu\varphi-\alpha\ln(\tilde{r}))}$ $\begin{array}[]{c}\alpha\{P_{3},D\}+\nu\{P_{3},L_{3}\}\\ -\left(2c_{1}x_{3}\cdot H\right)+2c_{2}x_{3}\end{array}$
7 $\frac{\tilde{\omega}}{F(\tilde{\omega}+\tilde{r}^{2})+G(\tilde{\omega}-\tilde{r}^{2})}$ $\frac{\tilde{F}(\tilde{\omega}+\tilde{r}^{2})+\tilde{G}(\tilde{\omega}-\tilde{r}^{2})}{F(\tilde{\omega}+\tilde{r}^{2})+G(\tilde{\omega}-\tilde{r}^{2})}$ $\begin{array}[]{c}L_{3}^{2}+(P_{1}^{2}-P_{2}^{2})\\ +\left(g(F,G)\cdot H\right)-g(\tilde{F},\tilde{G})\end{array}$
8 $\begin{array}[]{c}\frac{\tilde{\omega}}{\sum_{m}c_{m}\Omega_{m}},\end{array}$ $\begin{array}[]{c}\frac{\sum_{m}\tilde{c}_{m}\Omega_{m}}{\sum_{m}c_{m}\Omega_{m}}\end{array}$ $\begin{array}[]{c}L_{3}^{2}+(P_{1}^{2}-P_{2}^{2})\\ -2\left(\frac{y\sum_{m}c_{m}\Omega_{m-1}}{\tilde{\omega}}\cdot H\right)\\ +2y\sum_{m}\tilde{c}_{m}\frac{\Omega_{m-1}}{\tilde{\omega}}\end{array}$
The next table, i.e., Table 3 includes inverse masses and potentials of superintegrable systems admitting three integrals of motion one of which is the shifts generator $P_{3}$. Three of them include arbitrary functions while the remaining one include arbitrary parameters whose standard number is equal to eight. Let us remind that the notation $\varphi$ is used for the Euler angle.
Finally, the last is Table 4 which includes four maximally superintegrable systems. All of them include arbitrary parameters whose number is varying from one to six.
The integrals of motion presented in Tables 2-4 are functionally independent. In additional there is a number integrals of motion which are functionally independent with the presented ones but are linearly independent. To find them it is sufficient to calculate the commutators of the presented operators with $P_{3}$. The number of linearly independent integrals of motion for the systems presented in Items 2, 3 and 6 of Table 3 is equal to equal to 3 while the remaining items represent the systems for which this number is equal to 2.
The number of linearly independent integrals of motion for the systems presented in Table 4 are equal to four (Item 1), five (Item 2), six (Item 3) and seven (Item 4).
No $f$ $V$ Integrals of motion
1 $\frac{1}{c_{1}x_{1}+c_{2}x_{2}+2c_{3}\tilde{r}^{2}+c_{4}}$ $\frac{c_{5}x_{1}+c_{6}x_{2}+2c_{7}\tilde{r}^{2}+c_{8}}{c_{1}x_{1}+c_{2}x_{2}+2c_{3}\tilde{r}^{2}+c_{4}}$ $\begin{array}[]{c}P_{1}^{2}-P_{2}^{2}\\ +\left((c_{2}x_{2}-c_{1}x_{1}+c_{3}(x_{2}^{2}-x_{1}^{2}))\cdot H\right)\\ +c_{5}x_{1}-c_{6}x_{2}+c_{7}(x_{1}^{2}-x_{2}^{2}),\\ P_{1}P_{2}\\ -\frac{1}{2}\left(((2c_{3}+c_{2})x_{1}+c_{1}x_{2})\cdot H\right)\\ +\frac{1}{2}((2c_{7}+c_{6})x_{1}+c_{5}x_{2})\end{array}$
2 $\frac{1}{c_{1}x_{1}+F(x_{2})}$ $\frac{c_{2}x_{1}+\tilde{F}(x_{2})}{c_{1}x_{1}+F(x_{2})}$ $\begin{array}[]{c}P_{1}^{2}-P_{2}^{2}+c_{1}x_{1}-\tilde{F}(x_{2})\\ +\left((F(x_{2})-c_{1}x_{1})\cdot H\right),\\ \{P_{3},L_{2}\}-\frac{c_{1}}{2}(x_{3}^{2}\cdot H)+\frac{1}{2}c_{2}x_{3}^{2}\end{array}$
3 $\frac{x_{2}^{2}}{x_{2}^{2}F(x_{1})+c_{1}}$ $\frac{x_{2}^{2}\tilde{F}(x_{1})+c_{2}}{x_{2}^{2}F(x_{1})+c_{1}}$ $\begin{array}[]{l}L_{1}^{2}-\left(\frac{c_{1}x_{3}^{2}}{x_{2}^{2}}\cdot H\right)+\frac{c_{2}x_{3}^{2}}{x_{2}^{2}},\\ P_{1}^{2}-P_{2}^{2}+\left((\frac{c_{1}}{x_{2}^{2}}-F(x_{1}))\cdot H\right)\\ -\frac{c_{2}}{x_{2}^{2}}+\tilde{F}(x_{1})\end{array}$
6 $\frac{\tilde{r}^{2}}{c_{1}\tilde{r}^{2}+F(\varphi)}$ $\frac{c_{2}\tilde{r}^{2}+G(\varphi)}{c_{1}\tilde{r}^{2}+F(\varphi)}$ $\begin{array}[]{c}L_{3}^{2}-(F(\varphi)\cdot H)-G(\varphi),\\ D^{2}-c_{1}(r^{2}\cdot H)+c_{2}r^{2}\end{array}$
7 $\frac{\tilde{r}^{2}}{c_{1}(\mu\ln(\tilde{r})-\nu\varphi)+c_{2}\tilde{r}^{2}+c_{3}}$ $\frac{c_{4}(\mu\ln(\tilde{r})-\nu\varphi)+c_{5}\tilde{r}^{2}+c_{6}}{c_{1}(\mu\ln(\tilde{r})-\nu\varphi)+c_{2}\tilde{r}^{2}+c_{3}}$ $\begin{array}[]{c}\mu P_{3}L_{3}+\nu\{P_{3},D\}\\ -\left(2c_{2}\nu x_{3}\cdot H\right)+2c_{5}\nu x_{3},\\ L_{3}^{2}+2(c_{1}\nu\varphi\cdot H)-2c_{4}\nu\varphi\end{array}$
No $f$ $V$ Integrals of motion
1 $\frac{1}{x_{1}}$ $\frac{c_{1}x_{2}}{x_{1}}$ $\begin{array}[]{l}P_{1}^{2}-P_{2}^{2}-c_{1}x_{2}-\left(x_{1}\cdot H\right),\\ 2P_{1}P_{2}+c_{1}x_{1}-(x_{2})\cdot H),\\ \{P_{3},L_{2}\}-\frac{1}{2}(x_{3}^{2}\cdot H),\ \{P_{3},L_{1}\}-\frac{c_{1}x_{3}^{2}}{2}\end{array}$
2 $\frac{x_{1}^{2}x_{2}^{2}}{c_{1}x_{1}^{2}+c_{2}x_{2}^{2}+c_{3}x_{1}^{2}x_{2}^{2}}$ $\frac{c_{4}x_{1}^{2}+c_{5}x_{2}^{2}+c_{6}x_{1}^{2}x_{2}^{2}}{c_{1}x_{1}^{2}+c_{2}x_{2}^{2}+c_{3}x_{1}^{2}x_{2}^{2}}$ $\begin{array}[]{l}L_{3}^{2}-\left(\frac{c_{1}x_{1}^{4}+c_{2}x_{2}^{4}}{x_{1}^{2}x_{2}^{2}}\cdot H\right)+\frac{c_{5}x_{1}^{4}+c_{6}x_{2}^{4}}{x_{1}^{2}x_{2}^{2}},\\ L_{1}^{2}-\left(\frac{c_{1}x_{3}^{2}}{x_{2}^{2}}\cdot H\right)+\frac{c_{4}x_{3}^{2}}{x_{2}^{2}},\\ D^{2}-\left(c_{3}x_{3}r^{2}\cdot H\right)+c_{6}x_{3}r^{2}\end{array}$
3 $\frac{\tilde{r}x_{1}^{2}}{c_{1}x_{2}+c_{2}\tilde{r}+c_{3}\tilde{r}x_{1}^{2}}$ $\frac{c_{4}x_{2}+c_{5}\tilde{r}+c_{6}\tilde{r}x_{1}^{2}}{c_{1}x_{2}+c_{2}\tilde{r}+c_{3}\tilde{r}x_{1}^{2}}$ $\begin{array}[]{l}L_{3}^{2}-\left(\frac{c_{1}\tilde{r}x_{2}+c_{2}x_{2}^{2}}{x_{1}^{2}}\cdot H\right)-\frac{c_{4}\tilde{r}x_{2}+c_{5}x_{2}^{2}}{x_{1}^{2}},\\ D^{2}-\left(c_{2}r^{2}\cdot H\right)+c_{5}r^{2},\\ \{L_{2},L_{3}\}-\frac{2c_{5}x_{3}\tilde{r}+c_{4}x_{3}(x_{1}^{2}+2x_{2}^{2})}{x_{1}^{2}\tilde{r}}\\ +\left(\frac{2c_{2}x_{3}\tilde{r}+c_{1}x_{3}(x_{1}^{2}+2x_{2}^{2})}{x_{1}^{2}\tilde{r}}\cdot H\right)\end{array}$
4 $\frac{x_{1}^{2}}{x_{2}x_{1}^{2}+c_{1}}$ $\frac{c_{2}x_{2}x_{1}^{2}+c_{3}+c_{4}x_{1}^{2}}{x_{2}x_{1}^{2}+c_{1}}$ $\begin{array}[]{l}L_{2}^{2}-\left(\frac{c_{1}x_{3}^{2}}{x_{1}^{2}}\cdot H\right)+\frac{c_{3}x_{3}^{2}}{x_{1}^{2}},\\ \{P_{2},D\}-\frac{1}{2}\left((x_{1}^{2}+4x_{2}^{2}+x_{3}^{2})\cdot H\right)\\ +\frac{c_{2}}{2}(x_{1}^{2}+4x_{2}^{2}+x_{3}^{2})+2c_{4}x_{2},\\ \{P_{3},L_{1}\}+\left(\frac{c_{1}x_{3}^{2}}{2}\cdot H\right)-\frac{1}{2}c_{3}x_{3}^{2}\end{array}$
@@ 5 Discussion
We classify inequivalent quantum mechanical systems with position dependent masses which admit second order integrals of motion and one out of two one parametric Lie groups, i.e., dilatation and shift ones. The total number of such systems is equal to twenty seven, including nine ones invariant w.r.t. the dilatation transformations and eighteen shift invariant systems. The classification results are presented in four tables and include eleven integrable, ten superintegrable and six maximally superintegrable systems. Let us note that this statement is a bit conventional since the obtained systems are defined up to arbitrary functions and (or) arbitrary parameters, and in fact we enumerate the subclasses of the systems under study including the infinite sets of them. One more note is that the systems represented in Item 5 of Table 1 and Item 8 of Table 2 form the subclasses of the systems represented in Item 4 of Table 1 and Item 7 of Table 2 correspondingly.
The presented results can be treated as the next step in our programm of the complete classification of the 3d PDM systems admitting second order integrals of motion. Such systems possessing the symmetry w.r.t. three and two parametric Lie groups were classified in paper [42] and [45] respectively. The next step presupposes the classification of the systems admitting at least one parametric Lie symmetry group. In accordance with [24], up to equivalence it is sufficient to restrict ourselves to the two subclassess of the groups. The first of them includes rotations around the fixed axis, shifts along this axis and dilatations, we call these symmetries natural. The second subclass includes the so called combined transformations and includes superpositions of shifts and rotations, dilatations and rotations, and shift, rotations and conformal transformations.
The main result of the present paper consists in completing the classification of the integrable systems admitting the natural symmetries started in paper [47]. Thus to complete the classification of the 3d PDM systems admitting second order integrals of motion and a one parametric Lie symmetry group it is sufficient to classify the the systems admitting the combined symmetries mentioned in the above. This work is in progress.
One out of many stimulus to search for integrable and superintegrable hamiltonian systems consists in the fact that such systems as a rule are exactly solvable. This statement is an element of common knowledge for standard quantum mechanical systems. However, the same is true for systems with position dependent masses. In particular, it is the case for the rotationally invariant superintegrable PDM systems. It was shown in paper [43] that all such systems are exactly solvable. Moreover, in addition to the superintegrability, they are shape invariant with respect to the Darboux transform and can be effectively solved using the tools of supersymmetric quantum mechanics. Exact solvability of the PDM systems with extended Lie symmetries was prowed in paper [49].
The testing for exact solvability of superintegrable PDM systems lies out of frame of the present paper. However, it is possible to declare that at least some of the found systems do are separable and exactly solvable. In particular, it is the case for the system presented in Item 1 of Table 4. The study of the exact solvability aspects of the classified systems is one more problem which we plane to consider in the future research.
@@
6 Appendix
In our previous papers, e.g., in [46] we following form of hamiltonian (2):
$H=p_{a}fp_{a}+\hat{V}.$ (181)
$\hat{V}=V-V^{k}$ (182)
where
and $\Delta$ is the Laplace operator.
Formula (183) represents an example of kinematical potentials which can be reduced to zero by the rearranging the ambiguity parameters. One of the typical properties of such potentials is their independence on coupling constants.
The form (181) of the Hamiltonian is more nice than (3), however the corresponding potential is not necessary free of the kinematical part. On the other hand Hamiltonian in form (3) does not include such part. In addition, the latter form is more convenient for using the Stäckel transform with can be made by the simultaneous multiplication of this operator by the square roots of the potential from the l.h.s. and r.h.s. And it is why we use just realization (3) in the present paper.